SIPB IAP 2006 Classes Guide

http://sipb.mit.edu/iap/

Your guide to SIPB’s IAP 2006 class offerings

This small three-page excuse for a booklet you hold in your hands is your handy little reference to all the technical (or not-so-technical) classes you could ever hope to cram into your mind over the course of a month. Whether it’s the traditional crash courses in stock languages (Java/C/C++), a little dabbling in material for classes you plan to take, safeguarding surfing privacy, or creating armor... SIPB has the class for you.

If you have any questions, comments, concerns, or fan letters, please send them our way!

sipb-iap-sucker@mit.edu SIPB / W20-557 / MIT

(617) 253-7788 / x3-7788

See you in the classroom,

Mike Shaw SIPB IAP sucker

Contents

Linux/Solaris 2 Make Athena Work for You . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Athena Tips and Tricks, Theory and Practice (Beginner/Intermediate) . . . . . . . . . . . . . . 2 Gentoo Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Introduction to Gentoo Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Advanced Gentoo Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Solaris Internals: A Quick Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Other 3 A Preview of 6.033: Unix, DNS, and Cryptography . . . . . . . . . . . . . . . . . . . . . . . . . 3 Creating a Firefox Extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Wikis, Blogs, and Photo Galleries Made Easy: How to Use scripts.mit.edu to Create a Super-

charged Website . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 scripts.mit.edu and sql.mit.edu: User Forum and Q&A Session . . . . . . . . . . . . . . . . . . . 4 Information Security Lecture Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Leading Change: Charting the Course for the Department of the Navy’s IM/IT Transformation 4 Law Enforcement and Corporate IT - Convergence and Conflict . . . . . . . . . . . . . . . 4 Anonymous Communications for Crypto Geeks, the U.S. Department of Defense, and You . 5

Inessential Matlab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Athena and Your Laptop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 CosmosWorks Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 SolidWorks Seminar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

Other other 8 How to Make Chainmail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Good Experience Live 2005 Conference Sessions . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

Programming 8

The Crash Course Series: C, C++, Java . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Crash Course in Java . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Crash Course in C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Crash Course in C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Caffeinated Crash Course in C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Caffeinated Crash Course in C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Programming in PostScript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Shell Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 XML and XSLT in Web Development . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Web 2.0 development using Ruby on Rails . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Intro Programming with “Processing” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Network Programming with Python . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Programming Perl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 The Functional Programming Language Haskell . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Python Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Linux/Solaris

Make Athena Work for You

Janet Ryu, sipb-iap-athena@mit.edu Tue Jan 10, Thu Jan 12, 06-07:00pm, 1-115 No enrollment limit, no advance sign up. Participants welcome at individual sessions (series). No prereqs.

Do you dread having to use Athena? Is computer stuff “not your thing”? Since when did printing become so difficult? Come to this class and making “computing at MIT” a lot more enjoyable.

This class has two sessions which cover different material. The first class will cover topics that will make using Athena much easier, such as basic unix commands and avoiding common mishaps. The second session will cover ways to make switching between personal computer and athena much easier, including different ways to transfer files, print files, and access files.

Athena Tips and Tricks, Theory and Practice (Beginner/Intermediate)

Chaim Kutnicki, sipb-iap-tips@mit.edu Wed Jan 11, 06-08:00pm, 1-150 No enrollment limit, no advance sign up. Single session event.

Come and learn about the inner workings of Athena and how you can make Athena work for you. We’ll discuss basic concepts such as Kerberos, Hesiod and MIT’s mail system and what you can do to customize your computing experience to fit your needs. If time permits, we will demonstrate converting a Windows PC to a dual-boot Windows/Linux-Athena system.

Gentoo Linux

Introduction to Gentoo Linux

Rajiv Manglani, sipb-iap-gentoo@mit.edu Tue Jan 17, 08-10:00pm, 2-105 No enrollment limit, no advance sign up. Single session event. Prereq: Linux or UNIX experience.

Gentoo Linux is a versatile and fast distribution for x86, PowerPC, Sparc, Alpha, AMD64, IA64, and MIPS systems that is geared towards Linux power users. Unlike other distros, Gentoo Linux has an advanced package management system called Portage. Portage is a true ports system in the tradition of BSD ports, but is Python-based and sports a number of advanced features including dependencies, fine-grained package management, “fake” (OpenBSD-style) installs, path sandboxing, safe unmerging, system profiles, virtual packages, config file management, and more. A completed Gentoo Linux installation is customized and optimized for your specific needs. We will have an overview presentation and a demo of a running Gentoo system.

Advanced Gentoo Linux

Rajiv Manglani, sipb-iap-gentoo@mit.edu Thur Jan 19, 08-10:00pm, 2-105 No enrollment limit, no advance sign up. Single session event. Prereq: Some Gentoo experience.

This session will cover advanced usage of Gentoo Linux, a versatile and fast distribution for x86, PowerPC, Sparc, Alpha, AMD64, IA64, and MIPS systems that is geared towards Linux power users.

Details of Portage, the Gentoo package management system, will be presented. We will discuss creation of ebuild scripts, usage of various system tools such as qpkg and etcat. We will also have a detailed look into the init scripts system, and the various settings in make.conf. Current Gentoo users are encouraged to attend with their questions and anecdotes.

Solaris Internals: A Quick Overview

Bill Sommerfeld, sipb-iap-solaris@mit.edu Wed Jan 11, 08-10:00pm, 1-150 No enrollment limit, no advance sign up. Single session event. Prereq: Interest in OS internals; some prior exposure to unix.

A lightning fast tour through the Solaris operating system, touching on system architecture, boot, memory management, filesystems, networking, observability, and its transition towards an open development model.

Other

A Preview of 6.033: Unix, DNS, and Cryptography

Jennifer Tu, Tim Abbott, Kevin Chen, sipb-iap-6033@mit.edu Mon Jan 9, Mon Jan 23, 30, 06-07:00pm, 1-115 No enrollment limit, no advance sign up. Participants welcome at individual sessions (series).

Are you taking 6.033 this spring? Or next spring? Or just want to learn something new? We’ll talk about three topics covered in 6.033 – Unix, DNS, and cryptography. As a quick disclaimer, this IAP series only offers a small sampling of what you’ll experience in 6.033. Hopefully, it will help you understand the practical aspects and basics of how things work, so you can concentrate on the concepts and reasonings when you take the course.

Creating a Firefox Extension

Jeff Walden, sipb-iap-firefox@mit.edu Thu Jan 19, Tue Jan 24, Thu Jan 26, Tue Jan 31, Thu Feb 2, 07-09:00pm, 1-190 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: Proficiency with CSS/JavaScript/XML and some programming exp.

This class will teach you how to create a moderately complex Firefox extension which demonstrates most of the technologies used to create many extensions as well as Firefox itself. You will learn how to: write an extension’s UI, write a simple XPCOM component using JavaScript, use JavaScript to drive UI interactions, create a custom widget using XBL, and how to package an extension’s code to make an installable extension.

Wikis, Blogs, and Photo Galleries Made Easy: How to Use scripts.mit.edu to Create a Supercharged Website

Jeff Arnold, Joe Presbrey, sipb-scripts@mit.edu Fri Jan 20, 06-07:00pm, 32-124 No enrollment limit, no advance sign up. Single session event.

A new service now makes it possible to easily create an MIT website with one or more personal wikis, blogs, and/or photo galleries. We will introduce the service, demonstrate it, and answer any questions.

scripts.mit.edu and sql.mit.edu: User Forum and Q&A Session

Jeff Arnold, Joe Presbrey, sipb-scripts@mit.edu Fri Jan 27, 06-07:00pm, 32-124 No enrollment limit, no advance sign up. Single session event. All scripts.mit.edu and sql.mit.edu users are invited to come to this forum hosted by the creators of the service. Check out how others use the service and ask any questions you might have.

Information Security Lecture Series

Sherri Davidoff, sipb-iap-lectures@mit.edu Thu Jan 19, Thu Jan 26, Thu Feb 2, 07:30-09:30pm, 4-237

Leading Change: Charting the Course for the Department of the Navy’s IM/IT Transformation

David Wennergren, CIO of the Department of the Navy Thu Jan 19, 7:30-9:00PM, 4-237

The Department of the Navy is in the midst of a dramatic transformation to leverage the power of the “information age.” Mr. David Wennergren, the Chief Information Officer for the Department of the Navy will discuss the significant accomplishments of the Navy - Marine Corps team in achieving network-centric operations and knowledge dominance. He will also discuss the importance of successfully addressing change management as a part of this digital revolution, offering insights on challenges and lessons learned in dealing with cultural change and aligning the work of large, decentralized organizations.

Law Enforcement and Corporate IT - Convergence and Conflict

Detective Lieutenant John J. McLean, NEMLEC Computer Crime Unit (serving 46 departments), Medford Police Department Thu Jan 26, 07:30-09:00pm, 4-237

Law enforcement must be technically and tactically prepared to respond to criminal activity in a legally sound and efficient manner. Corporations must protect their assets with solid business continuity plans and proper protocols and notifications. These goals can be in direct conflict. This lecture will introduce and examine the interplay between the so-called L.E. model (identifying the culprits and seizing evidence), and the corporate response of business protection and continuity. Emphasis on practical security measures and reasonable enforcement approaches that are based on the principle of “community policing.”

Anonymous Communications for Crypto Geeks, the U.S. Department of Defense, and You

Nick Mathewson and Roger Dingledine, developers of Tor Thu Feb 2, 07:30-09:30pm, 4-237

Anonymity for the masses is approaching. We’re developing Tor (tor.eff.org), a tool for strong anonymity and privacy on the Internet. Funded by the DoD and the EFF, Tor is a distributed overlay network of privacy servers around the world. First half: Introduction to the theory and designs that make anonymity networks work, and a discussion of the technical choices from the earliest designs of the 1980s to today. Second half: Experiences deploying anonymity, including backing, adoption, law enforcement issues, abuse, anti-abuse measures, usability, co-existing with other systems, censorious governments, and creating incentives for people to help privacy. We’ll also discuss threats to anonymous networks from legal, social, and technical avenues.

Inessential Matlab

Chris Porter, sipb-iap-matlab@mit.edu Wed Jan 25, 05:30-07:30pm, 4-231 No enrollment limit, no advance sign up. Single session event.

Introductory class on Matlab.

Athena and Your Laptop

Adam Seering, sipb-iap-laptop@mit.edu Tue Jan 17, Thu Jan 19, Tue Jan 24, Thu Jan 26, 01:00-02:30pm, 37-212 No enrollment limit, no advance sign-up required. Recommended: Have a computer on campus, laptop or desktop.

Tired of having to walk all the way to an Athena cluster to use Athena-specific services? Well, you don’t have to! Learn to use your own personal computer (either a laptop or a desktop) to take full advantage of MITNet. Topics covered include using an e-mail program (Thunderbird, Eudora, Apple Mail) instead of WebMail, printing to Athena printers, accessing your Athena home directory and other AFS lockers, and running software such as Matlab and Maple. Both Mac and Windows PC setups will be discussed; if there is interest, Linux setups will be discussed too. Time permitting, we’ll cover any requested topics as well. If you have a laptop, bring it!

CosmosWorks Seminar

Alex Slocum, Jr., sipb-iap-cosmosworks@mit.edu Thu Jan 12, 06-09:00pm, 4-231 No enrollment limit, no advance sign up. Single session event.

From Installation of Cosmos Works from the MIT Server, to CosmosWorks studies for individual parts as well as assemblies of parts, come get your CosmosWorks questions answered. Basic CosmosMotion and Cosmos FloWorks will also be covered if enough interest is shown. (Note: Laptops WILL NOT be provided). calendar

6 calendar

7

SolidWorks Seminar

Alex Slocum, Jr., sipb-iap-solidworks@mit.edu Wed Jan 11, 06-09:00pm, 4-231 No enrollment limit, no advance sign up. Single session event.

From Installation of SolidWorks from the MIT Server, to advanced assemblies, come get your SolidWorks questions answered (Note: Laptops WILL NOT be provided).

Other other

How to Make Chainmail

Cyrus Eyster, sipb-iap-chainmail@mit.edu Mon Jan 23, 09-10:00pm, 1-132 No enrollment limit, no advance sign up. Single session event.

Learn to make chainmail! This flexible armor made from metal rings was used on battlefields across the world for over a thousand years. This class will teach participants the basic weave used in Europe in the Middle Ages. Tools and materials will be provided.

Good Experience Live 2005 Conference Sessions

Rajiv Manglani, sipb-iap-gel@mit.edu Mon Jan 30, 07-08:30pm, 1-150 No enrollment limit, no advance sign up. Single session event. Prereq: interest in user experience.

Gel (“Good Experience Live” is a conference and community exploring good experience in all its forms – in business, art, society, technology, and life. Join us as we watch sessions from the 2005 conference DVD: Barry Schwartz, author of “The Paradox of Choice” Bob Mankoff, The New Yorker cartoon editor; Jimmy Wales, Wikipedia founder; Charlie Todd, urban prankster; Theo Jansen, strandbeest; and The Flying Karamazov Brothers! If you plan on attending please RSVP to sipb-iap-gel@mit.edu.

Programming

The Crash Course Series: C, C++, Java

Crash Course in Java

Ricky Ramirez, sipb-iap-java@mit.edu Mon Jan 23 thru Thu Jan 26, 04-06:00pm, 2-105 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series).

This class is an introduction to the Java programming language. The first half will cover basics of the language, including Java development environments on Athena and language syntax. The second half will examine the Java run-time library, including standard container classes, file I/O, and graphical user interfaces using Swing.

Crash Course in C

Michael Shaw, sipb-iap-c@mit.edu Tue-Thu, Jan 17, 19, 24, 26, 31-1, 06-08:00pm, 1-150 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: Minimal programming experience.

Interested in C, but not enough time for 1.00 or 10.001? This course will cover the fundamentals of programming in C, including basic control structures, functions, pointers and arrays, and an introduction to the C standard library.

Crash Course in C++

Nelson Elhagae, sipb-iap-seeplusplus@mit.edu Tue, Thu, Jan 10, 12, 17, 19, 24, 26, 03-05:00pm, 37-212 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: Familiarity with programming, including object-orientation.

This class will cover a crash course in C++ for coders who are already basically proficient with code and object-oriented programming. We will cover C++ syntax and concepts, including objects and namespace, and get into a quick intro to the standard template library. No prior experience with C is assumed.

Caffeinated Crash Course in C++

Mike Rolish, sipb-iap-caffeinatedplusplus@mit.edu Mon Jan 23, 01-04:00pm, 1-190 No enrollment limit, no advance sign up. Single session event. Prereq: Some programming experience.

A three-hour introduction to the C++ programming language. Basic syntax, data types, and control flow. Object basics. Dynamic memory management. Polymorphism and templates. Standard library. Useful tools.

Caffeinated Crash Course in C

Steve Summit, sipb-iap-cccc@mit.edu Wed Jan 25, 4-7pm and optional 8-10pm, 1-150

C’s influence is deeply pervasive in today’s software systems, and in the many currently-popular programming languages derived from C. In fact, C plays a role somewhat similar to the one once played by assembly language: even if you don’t do any actual day-to-day C programming, knowing something about C can be a huge help in better understanding the other systems and languages you are working with.

This class will attempt to cover the entire C programming language in 3 hours, in the grand tradition of SIPB’s IAP caffeinated crash courses. Prior programming experience, preferably in a C-like language (C++, Java, Perl, etc.) is very definitely assumed. The focus will be on understanding why the language is as it is, and on the implications for existing programs written in C or for work in newer programming languages derived from C.

There will be a special bonus 2-hour section following the first half, for those whose brains have not yet leaked out of their ears. The instructor is an MIT alumnus and maintainer of the C FAQ list on the ’net.

Programming in PostScript

Bayard Wenzel, sipb-iap-postscript@mit.edu Mon Jan 23, Wed Jan 25, Mon Jan 30, Wed Feb 1, 11am-12:00pm, 26-204 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: Familiarity with programming languages.

PostScript is the standard page description language, and the language which PDF, the standard document interchange format, is based upon. This class provides an introduction to PostScript, how to write both PostScript programs and programs that write PostScript, and an overview of the structure and construction of PDF files.

Shell Scripts

Jacob Morzinski, sipb-iap-shell@mit.edu Mon Jan 23, 02-03:30pm, 2-132 No enrollment limit, no advance sign up. Single session event. Prereq: Some experience with Unix and basic tools (ls, echo, cat).

Shell scripts are the common method of automating command-prompt tasks on Unix systems. This class will introduce the shell (the command-prompt), explain how the shell reacts to your typing, and present the fundamentals of shell programming.

XML and XSLT in Web Development

William Reilly, sipb-iap-xml@mit.edu Cosponsor: Libraries Tue Jan 17, Wed Jan 18, Tue Jan 24, 05:30-07:30pm, 4-231 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: Familiarity with XML, some exposure to XSLT transformations.

Introduction to the use of fundamental XML technologies in website development, in particular XSLT (and XPath). Attention also to XML Schemas and DTDs. Demonstration and investigation of the web-based XML publishing framework Cocoon. Examples of use of specific markup languages (applications of XML) relevant to website development (XHTML, SVG, DocBook, RSS).

Web 2.0 development using Ruby on Rails

Matt DeBergalis, sipb-iap-ruby@mit.edu Wed Feb 1, Thu Feb 2, 07-09:00pm, 4-370 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: Minimal comfort with HTML and scripting.

Hosted web services have garnered considerable attention in the past year as an attractive alternative to traditional desktop software. We’ll walk through how Ruby on Rails – a new web application framework – lets one quickly build and deploy a modern “Web 2.0” application that takes advantage of AJAX, REST, SOAP, and much much more, using code powering ActBlue.com as a working example. No Ruby or previous web development experience necessary.

Intro Programming with “Processing”

J.D. Zamfirescu, sipb-iap-processing@mit.edu Tue Jan 24, Thu Jan 26, 08-10:00pm, 1-115 Enrollment limit 40, no advance sign up. Repeating event. Participants welcome at any session.

Bored with “Hello World?” Learn how to program with a graphics- oriented language: Processing. See the Processing website (www.processing.org) for examples of the type of code (and designs!) we’ll be producing. Please feel free to bring a laptop!

Network Programming with Python

J.D. Zamfirescu, sipb-iap-network@mit.edu Tue Jan 24, Thu Jan 26, 04-06:00pm, 1-115 Enrollment limit 20, no advance sign up. Repeating event. Participants welcome at any session. Prereq: Basic knowledge of Python...see tutorial at www.python.org

Learn about sockets and network programming using the Python language! Over the course of two hours, we’ll talk about Internet addressing, ports, and the server/client architecture; we’ll finish off the class by writing a brief web server. Please feel free to bring a laptop!

Programming Perl

Alex Rolfe, sipb-iap-perl@mit.edu Tue Jan 17, Thu Jan 19, Tue Jan 24, Thu Jan 26, 08-10:00pm, 1-150 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: some programming experience.

Introduction to programming in Perl: syntax, flow control, I/O, regular expressions, data structures, objects, and some CGI programming.

The Functional Programming Language Haskell

David Glasser, sipb-iap-haskell@mit.edu Mon Jan 23 thru Thu Jan 26, 02-03:00pm, 2-105 No enrollment limit, no advance sign up. Participants requested to attend all sessions (non-series). Prereq: 6.001 or other programming experience recommended.

This is an introduction to the programming language Haskell. Haskell is a functional programming language (like Scheme), with types (like ML), and lazy evaluation (unlike almost any other language). We will get introduced to the language, create a library for image manipulation, and touch on advanced topics such as monadic programming.

Python Programming

Giles Hall, sipb-iap-progpython@mit.edu Tue Jan 10, Thu Jan 12, Tue Jan 17, 07-10:00pm, 1-190 No enrollment limit, no advance sign up. Participants welcome at individual sessions (series).

Tuesday January 10: Introduction to Python Programming: Python is easy to learn, the basics of the scripting language are taught in the first class. Thursday January 12: Object Oriented Python Programming: Built to be an object-oriented language from the ground up, the second class shows off the ease and power of python objects. Tuesday January 17: Advanced Python Techniques: Learn to speed-up your code by extending the python interpreter using its C API. meep?

13

Monday  Tuesday  Wednesday  Thursday  Friday
9  10  11  12  13
6.033 Sampler  CCC++  CCC++
6-7pm, 1-115  3-5pm, 37-212  3-5pm, 37-212
Make Athena Work for you  Athena Tips and Tricks  Make Athena Work for you
6-7pm, 1-115  6-8pm, 1-150  6-7pm, 1-115
SolidWorks  CosmosWorks
6-9pm, 4-231  6-9pm, 4-231
Python  Solaris Internals  Python
7-10pm, 1-190  8-10pm, 1-150  7-10pm, 1-190
16  17  18  19  20
Laptop Athena  Laptop Athena
1-2:30pm, 37-212  1-2:30pm, 37-212
CCC++  CCC++
3-5pm, 37-212  3-5pm, 37-212
CCC  CCC  Wikis and Blogs
6-8pm, 1-150  6-8pm, 1-150  6-7pm, 32-124
Python  Firefox Extensions
7-10pm, 1-190  7-9pm, 1-190
XML/XSLT in webdev  XML/XSLT in webdev  Lecture Series - Navy
5:30-7:30pm, 4-231  5:30-7:30pm, 4-231  7:30-9pm, 4-237
Gentoo (Intro)  Gentoo (Advanced)
8-10pm, 2-105  8-10pm, 2-105
Perl  Perl
8-10pm, 1-150  8-10pm, 1-150
23  24  25  26  27
6.033 Sampler  Laptop Athena  Laptop Athena
6-7pm, 1-115  1-2:30pm, 37-212  1-2:30pm, 37-212
Haskell  Haskell  Haskell  Haskell
2-3pm, 2-105  2-3pm, 2-105  2-3pm, 2-105  2-3pm, 2-105
Postscript  Python netprog  Postscript  Python netprog
11am-12pm, 26-204  4-6pm, 1-115  11am-12pm, 26-204  4-6pm, 1-115
CCCC++  CCC++  CCCC  CCC++
1-4pm, 1-190  3-5pm, 37-212  4-7pm, 1-150  3-5pm, 37-212
Shell Scripts  CCC  Matlab  CCC
2-3pm, 2-132  6-8pm, 1-150  5:30-7:30pm, 4-231  6-8pm, 1-150
CCJava  CCJava  CCJava  CCJava
4-6pm, 2-105  4-6pm, 2-105  4-6pm, 2-105  4-6pm, 2-105
XML/XSLT in webdev  Lecture Series - Medford Police
5:30-7:30pm, 4-231  7:30-9pm, 4-237
Chainmail  Firefox Extensions  Firefox Extensions  scripts user forum
9-10pm, 1-132  7-9pm, 1-190  7-9pm, 1-190  6-7pm, 32-124
Proc. programming  Proc. programming
8-10pm, 1-115  8-10pm, 1-115
Perl  Perl
8-10pm, 1-150  8-10pm, 1-150
30  31  1  2  3
Postscript  CCC  Postscript  CCC
11am-12pm, 26-204  6-8pm, 1-150  11am-12pm, 26-204  6-8pm, 1-150
6.033 Sampler  Firefox Extensions  Firefox Extensions
6-7pm, 1-115  7-9pm, 1-190  7-9pm, 1-190
Good XP Live  Lecture Series - tor
7-8:30pm, 1-150  7:30-9pm, 4-237
Ruby on Rails  Ruby on Rails
7-9pm, 4-370  7-9pm, 4-370